OpenStudy (anonymous):

1. x^1/3 * x^1/3 2. 4 ______________ (3x)^-2 3.Simplify. Express the product as a radical expression (x^3)^3/4 PLEASE HELP!!

OpenStudy (anonymous):

\[1. x^{1/3}\times x^{1/3}\]\[2.\frac{4}{3x^{-2}}\] but i still don't get the question o_o.. the product of 1 and 2?

OpenStudy (anonymous):

\[1. \huge x^{\frac{1}{3}+\frac{1}{3}}= x^{\frac{2}{3}}\] \[2. \huge \frac{4}{3x^{-2}}=\frac{4}{3}x^2\] \[\huge x^{\frac{2}{3}}\times \frac{4}{3}x^2= \frac{4}{3} x^{\frac{2}{3} + 2}\] \[\huge \frac{4}{3}x^{\frac{2}{3}+\frac{6}{3}}= \frac{4}{3}x^{\frac{8}{3}}\]